# Enhancement of non-Gaussianity and nonclassicality of pair coherent states with postselected von Neumann measurement

### v1

@@ Abstract
We investigate the effects of postselected von Neumann measurements on the nonclassical properties of pair coherent states (PCS). We calculated key quantum characteristics, such as squeezing, photon statistics, and entanglement between the two PCS modes. Our results demonstrate that postselected von Neumann measurements enhance both the non-Gaussianity and nonclassicality of PCS. These findings are validated by analyzing the scaled joint Wigner function across various system parameters. The theoretical optimization scheme offers an alternative approach for improving PCS-based quantum information efficiency and facilitates practical implementations in quantum technologies.
@@ I Introduction
The nonclassicality of quantum states plays a crucial role in quantum computation and information processing. Recent advancements in technology allow the manipulation of photons, making continuous-variable states of bosonic modes an excellent platform for quantum information processing. Among the variety of continuous-variable states (Lvovsky et al., 2020; Walschaers, 2021; Asavanant and Furusawa, 2024), two-mode Gaussian and non-Gaussian states such as entangled cat states (Wang et al., 2016), N00N states (Afek et al., 2010; Wang et al., 2011), squeezed states (Dao-ming, 2015; Riabinin et al., 2021), and pair coherent states (PCS) (Agarwal, 1986) have irreplaceable roles in numerous applications (Tara and Agarwal, 1994; Gábris and Agarwal, 2007; Zhou et al., 2010; Seshadreesan et al., 2015; Vahlbruch et al., 2016; Wang and Zhao, 2017; Acernese et al., 2019; Dat, 2023; Roman-Rodriguez et al., 2024). In two-mode nonclassical states, enhancing nonclassicality and entanglement properties can be achieved through photon addition and subtraction operations (Hong, 1999; Hong and Guang-can, 1999; Kitagawa et al., 2006; Biswas and Agarwal, 2007; Yuan et al., 2009; Truong et al., 2014; Hoai and Duc, 2016; Yuan et al., 2018; Lu, 2018; Duc et al., 2019, 2020; Truong et al., 2021; Chuong and Duc, 2023). However, operating with non-Gaussian states introduces additional computational complexity compared to Gaussian states. Beyond photon addition and subtraction (or their superpositions) (Parigi et al., 2007), quantum catalysis provides another feasible method for enhancing the nonclassicality of quantum states (Duc et al., 2020). Quantum catalysis enhances quantum properties without directly altering the photon count. This enhancement arises from the inherently quantum mechanical nature of the process (for further details, refer to (Sanaka et al., 2006), and references therein). However, quantum catalysis is not universally applicable to all states, necessitating alternative methods for optimizing nonclassical states. Specifically, exploring new state-optimization schemes is essential for improving the efficiency of multimode state-based applications–such as multimode Gaussian and non-Gaussian states–without relying on photon addition or subtraction operations.
@@
In this work, we investigate how the properties of PCS change after applying the theory of postselected measurements to one mode of PCS. We calculate the squeezing, quantum statistics, and entanglement characteristics of the measurement output state. Our results show that postselected von Neumann measurements could enhance the non-Gaussianity and nonclassicality of PCS, considering large weak values of the measured system observable and appropriate coupling strength parameters between the measured system and measuring device (MD). We confirm these findings by examining the phase-space distribution of our state, characterized by the scaled joint Wigner function.
We organized the paper as follows: In Sec. II, we present our theoretical model. Section III explores the effects of postselection on the squeezing properties of PCS, including quadrature and sum squeezing. In Sec. IV, we investigate the quantum statistics of postselected PCS under von Neumann measurements, analyzing the relationship between the $a$-mode and $b$-mode using the second-order cross-correlation (SOCC) and second-order correlation functions. Section V examines entanglement through measures such as the Hillery-Zubairy (HZ) correlation and Einstein–Podolsky–Rosen (EPR) correlation. In Sec. VI, we investigate the scaled joint Wigner function of our measurement output state and take comparison with the initial PCS case. Section VII explores the fidelity of the postselected PCS under von Neumann measurements. Finally, Sects. VIII and IX present a discussion and conclusion, respectively. Throughout this paper, we take $\hbar=1$.
@@ II MODEL SETUP
$|\gamma,\delta\rangle=\mathcal{N}_{n}\sum_{n=0}^{\infty}\frac{\gamma^{n+\delta/2}}{\sqrt{n!(n+\delta)!}}\left|n+\delta\right\rangle_{a}\left|n\right\rangle_{b}.$ (1)
Here, $\gamma$ is a complex number representing the amplitude and phase of the state, i.e., $\gamma=re^{i\varphi}$; $\delta$ is an integer denoting the photon number difference (PND) between the two modes; and $\mathcal{N}_{n}$ is a normalization coefficient defined by:
$\mathcal{N}_{n}$ $=\left(\sum_{n=0}^{\infty}\frac{|\gamma|^{2n+\delta}}{n!(n+\delta)!}\right)^{-\frac{1}{2}}=\frac{1}{\sqrt{I_{\delta}(2|\gamma|)}}.$ (2)
@@
Here, $a$ ($a^{\dagger}$) and $b$ ($b^{\dagger}$) are the annihilation (creation) operators associated with the two modes ($a$-mode and $b$-mode) of the PCS, respectively. The realization of PCS and the study of its characteristics have been achieved experimentally (Gertler et al., 2023). PCS are a type of two-mode non-Gaussian and nonclassical states (Hong and Guang-can, 1999), and their generations have been investigated both theoretically (Agarwal, 1986; Gou et al., 1996; Zheng, 2002; Obada and Khalil, 2006; Dong et al., 2008; Gerry et al., 2011) and experimentally (Gertler et al., 2023; Jeon et al., 2024). The sub-Poissonian statistics, correlations in number fluctuations, squeezing, and violations of Cauchy–Schwarz inequalities (CSI) have been studied extensively (Agarwal, 1988; Lee, 1990; Obada et al., 2020; Duc et al., 2019; Truong et al., 2021). As non-Gaussian entangled states, PCS offers advantages in performing quantum tasks such as quantum communication (Tara and Agarwal, 1994; GÁBRIS and AGARWAL, 2007; Albert et al., 2019) and quantum key distribution (Zhou et al., 2010; Wang and Zhao, 2017).
@@
To study the effects of postselected von Neumann measurement on the PCS, we consider the PCS as the MD (measuring device) and its polarization as the measured system. As illustrated in Fig. 1, we simplify the setup by restricting the interaction between the MD and the measured system to the $a$-mode. The interaction Hamiltonian $H_{int}$ is of the von Neumann type, expressed as:
@@
$\times$ $|\psi_{i}\rangle\otimes|\phi\rangle.$
where $\Gamma=gt/\sigma$ represents the ratio between $gt$ and $\sigma$, $\mathbb{I}$ is the $2\times 2$ identity matrix, and $D(\Gamma/2)=e^{\frac{\Gamma}{2}(\hat{a}^{\dagger}-\hat{a})}$ is the displacement operator acting on the $a$ mode of the PCS. It is important to note that the coupling strength parameter $\Gamma$ characterizes the measurement strength in this study. The measurement is referred to as weak when $\Gamma\text{$\ll 1$}$ and strong when $\Gamma\gg 1$. For this reason, $\Gamma$ can be considered the transaction parameter in our measurement model. The measurement process involves postselection. In this study, we assume the postselection of the system state to be $|\psi_{f}\rangle=|H\rangle$. After postselecting the system state, the final state of the MD is obtained, and its expression is given as
@@
with $P=\mathcal{N}_{n}^{2}e^{-\frac{\Gamma^{2}}{2}}\sum\limits_{n=0}^{\infty}\left[\frac{|\gamma|^{2n+\delta}}{n!(n+\delta)!}L_{n+\delta}^{(0)}\left(\Gamma^{2}\right)\right]$, and $\langle\sigma_{x}\rangle_{w}$ is the weak value of the system observable $\sigma_{x}$ defined as
@@
For simplify, we assume the weak value to be a real number by setting $\vartheta=0$. This weak value can be obtained in a measurement process with a postselection probability given by $P_{s}=|\langle\psi_{f}|\psi_{i}\rangle|^{2}=\cos^{2}(\frac{\alpha}{2})$. In calculating the normalization coefficient $\lambda$, the following expressions were used
@@ III.1 Quadrature squeezing
As we can see, the values of $Q_{i}$ are bounded as $Q_{i}\geq-\frac{1}{4}$, and the $i$-th component of the quadrature operators of the PCS is said to be squeezed if $-\frac{1}{4}\leq Q_{i}<0$. After some algebra, we can obtain the specific expressions for the quadrature squeezing parameter $Q_{i}$ of the final state $|\Psi\rangle$, expressed as:
@@
To provide a clearer analysis of quadrature squeezing, we rely on numerical calculations, with the results presented in Figs. 2 and 3. In Fig. 2, we plot $Q_{1,\Psi}$ as a function of the parameter $\gamma$ for various coupling strength parameters $\Gamma$, while the weak value $\langle\sigma_{x}\rangle_{w}=5.671$ is fixed, corresponding to $\alpha=8\pi/9$. As shown in Fig. 2, $Q_{1,\phi}$ ($\Gamma=0$) remains positive and increases with $\gamma$ [see the thick green curve]. Additionally, Fig. 2 demonstrates that $Q_{1,\Psi}$ for cases where $\Gamma\neq 0$ also consistently takes positive values. This observation confirms that after the postselected von Neumann measurement, the $F_{1}$ quadrature of the enhanced PCS $|\Psi\rangle$ still doesn’t exhibit squeezing, consistent with the behavior of the initial state.
@@
In Fig. 3, we present the numerical results for $Q_{2,\Psi}$, which corresponds to the quadrature $F_{2}$ of the enhanced PCS $|\Psi\rangle$. In Fig. 3(a), $Q_{2,\Psi}$ is plotted as a function of the parameter $\gamma$ for various coupling strength parameters $\Gamma$, with a fixed large weak value $\langle\sigma_{x}\rangle_{w}=5.671$, corresponding to $\alpha=8\pi/9$. Among these curves, the case $\Gamma=0$ (thick green curve) represents $Q_{2,\phi}$ for the initial state. As shown, $Q_{2,\phi}$ decreases with increasing $\gamma$ and converges to $-0.125$ across most parameter regions. Experimentally, the amount of squeezing is quantified in terms of decibels (dB) (Agarwal, 2013). This value of $-0.125$ corresponds to $9$ dB or $50\%$ squeezing. Notably, in the specific case where the PND is $\delta=0$, $Q_{2,\phi}$ is consistently less than zero. Thus, the initial PCS $|\phi\rangle$ with $\delta=0$ always exhibits squeezing along the $F_{2}$ quadrature.
@@
Figure 2: Quadrature squeezing $Q_{1,\Psi}$ as a function of the parameter $\gamma$. The thick green line corresponds to the interaction parameter $\Gamma=0$, representing the case where the state is the initial PCS $|\phi\rangle$. The solid gray line represents $\Gamma=0.3$, while the purple, blue, and red dashed lines correspond to $\Gamma=0.5$, $\Gamma=0.7$, and $\Gamma=1$, respectively. For these calculations, we take $\delta=0$, $\alpha=\frac{8\pi}{9}$ and $\varphi=0$.
@@
In Fig. 3(a), after the postselected von Neumann measurements ($\Gamma\neq 0$), we observe that the value of $Q_{2,\Psi}$ is lower than $Q_{1,\phi}$ for the weak interaction parameter $\Gamma=0.3$. This behavior persists over a larger range of the parameter $\gamma$, indicating that the squeezing characteristics of the quadrature $F_{2}$ are enhanced under postselected von Neumann measurements. However, as the interaction coupling strength parameter $\Gamma$ increases, the value of $Q_{2,\Psi}$ decreases further compared to the $\Gamma=0.3$ case. For $\Gamma=0.3$, the minimum value of $Q_{2,\Psi}$ reaches approximately $-0.172$ near $\gamma=10$. This corresponds to a squeezing level of $11$dB or $69\%$. Therefore, after performing postselected WMs, the quadrature squeezing along the $F_{2}$ direction of the field increases by approximately $19$% under suitable system parameters and weak values of the measured system observable. To further investigate the signal amplification effect of weak values for weak system signals, Fig. 3(b) presents $Q_{2,\Psi}$ as a function of the weak value angle $\alpha$ for different coupling strength parameters $\Gamma$ and a fixed $\gamma=10$. The numerical results in Fig. 3(b) indicate that the weak value amplifies the squeezing effects of $Q_{2,\Psi}$ when the interaction coupling strength parameter $\Gamma$ is small, highlighting the signal amplification capability of weak measurements. These results demonstrate that the squeezing effects of $Q_{2,\Psi}$ can be enhanced by postselected WMs.
@@
Figure 3: Quadrature squeezing $Q_{2,\Psi}$: (a) $Q_{2,\Psi}$ as a function of the parameter $\gamma$ for different coupling strength parameters $\Gamma$, with the weak value fixed at $\langle\sigma_{x}\rangle_{w}=5.671$ corresponding to $\alpha=8\pi/9$. (b) $Q_{2,\Psi}$ as a function of the weak value parameter $\alpha$ for different coupling strength parameters $\Gamma$, with the parameter $\gamma$ fixed at 10. Other parameters are the same as those in Fig. 2.

### v2

@@ Abstract
We investigate the effects of postselected von Neumann measurements on the nonclassical properties of pair coherent states (PCS). We calculated key quantum characteristics, such as squeezing, photon statistics, entanglement between the two PCS modes, and teleportation fidelity. Our results demonstrate that postselected von Neumann measurements enhance both the non-Gaussianity and nonclassicality of PCS. These findings are validated by analyzing the scaled joint Wigner function across various system parameters. We also notice that in weak measurement regimes, the teleportation fidelity via a PCS-based quantum channel under postselected von Neumann measurements could keep successful teleportation for anomalous weak values of the system obervable. The theoretical optimization scheme offers an alternative approach for improving PCS-based quantum information efficiency and facilitates practical implementations in quantum technologies.
@@ I Introduction
The nonclassicality of quantum states plays a crucial role in quantum computation and information processing. Recent advancements in technology allow the manipulation of photons, making continuous-variable states of bosonic modes an excellent platform for quantum information processing. Among the variety of continuous-variable states (Lvovsky et al., 2020; Walschaers, 2021; Asavanant and Furusawa, 2024), two-mode Gaussian and non-Gaussian states such as entangled cat states (Wang et al., 2016a), N00N states (Afek et al., 2010; Wang et al., 2011), squeezed states (Dao-ming, 2015; Riabinin et al., 2021), and pair coherent states (PCS) (Agarwal, 1986) have irreplaceable roles in numerous applications (Tara and Agarwal, 1994; Gábris and Agarwal, 2007; Zhou et al., 2010; Seshadreesan et al., 2015; Vahlbruch et al., 2016; Wang and Zhao, 2017; Acernese et al., 2019; Dat, 2023; Roman-Rodriguez et al., 2024). In two-mode nonclassical states, enhancing nonclassicality and entanglement properties can be achieved through photon addition and subtraction operations (Hong, 1999; Hong and Guang-can, 1999; Kitagawa et al., 2006; Biswas and Agarwal, 2007; Yuan et al., 2009; Truong et al., 2014; Hoai and Duc, 2016; Yuan et al., 2018; Lu, 2018; Duc et al., 2019, 2020; Truong et al., 2021; Chuong and Duc, 2023). However, operating with non-Gaussian states introduces additional computational complexity compared to Gaussian states. Beyond photon addition and subtraction (or their superpositions) (Parigi et al., 2007), quantum catalysis provides another feasible method for enhancing the nonclassicality of quantum states (Duc et al., 2020). Quantum catalysis enhances quantum properties without directly altering the photon count. This enhancement arises from the inherently quantum mechanical nature of the process (for further details, refer to (Sanaka et al., 2006), and references therein). However, quantum catalysis is not universally applicable to all states, necessitating alternative methods for optimizing nonclassical states. Specifically, exploring new state-optimization schemes is essential for improving the efficiency of multimode state-based applications–such as multimode Gaussian and non-Gaussian states–without relying on photon addition or subtraction operations.
@@
In this work, we investigate how the properties of PCS change after applying the theory of postselected measurements to one mode of PCS. We calculate the squeezing, quantum statistics, and entanglement characteristics of the measurement output state. Our results show that postselected von Neumann measurements could enhance the non-Gaussianity and nonclassicality of PCS, considering large weak values of the measured system observable and appropriate coupling strength parameters between the measured system and measuring device (MD). We confirm these findings by examining the phase-space distribution of our state, characterized by the scaled joint Wigner function. We also check the teleportation fidelity with postselected von Neumann measurements to the PCS and the numerical outcomes indicate that in weak measurement regimes the quantum teleportation task based on PCS-based quantum channel is still could keep its validity for larger weak values of measured system observable.
We organized the paper as follows: In Sec. II, we present our theoretical model. Section III explores the effects of postselection on the squeezing properties of PCS, including quadrature and sum squeezing. In Sec. IV, we investigate the quantum statistics of postselected PCS under von Neumann measurements, analyzing the relationship between the $a$ mode and $b$ mode using the second-order cross-correlation (SOCC) and second-order correlation functions. Section V examines entanglement through measures such as the Hillery-Zubairy (HZ) correlation and Einstein–Podolsky–Rosen (EPR) correlation. In Sec. VI, we investigate the scaled joint Wigner function of our measurement output state and take comparison with the initial PCS case. Section VII we present the state distance between initial and final MD states, and give the details of the teleportation fidelity of a coherent state transmitted through a PCS-based quantum channel under postselected von Neumann measurements. Finally, Sects. VIII and IX present a discussion and conclusion, respectively. Throughout this paper, we take $\hbar=1$.
@@ II MODEL SETUP
$|\gamma,\delta\rangle=\mathcal{N}_{\delta}\sum_{n=0}^{\infty}\frac{\gamma^{n+\delta/2}}{\sqrt{n!(n+\delta)!}}\left|n+\delta\right\rangle_{a}\left|n\right\rangle_{b}.$ (1)
Here, $\gamma$ is a complex number representing the amplitude and phase of the state, $\delta$ is an integer denoting the photon number difference (PND) between the two modes; and $\mathcal{N}_{\delta}$ is a normalization coefficient defined by:
$\mathcal{N}_{\delta}$ $=\left(\sum_{n=0}^{\infty}\frac{|\gamma|^{2n+\delta}}{n!(n+\delta)!}\right)^{-\frac{1}{2}}=\frac{1}{\sqrt{I_{\delta}(2|\gamma|)}}.$ (2)
@@
Here, $a$ ($a^{\dagger}$) and $b$ ($b^{\dagger}$) are the annihilation (creation) operators associated with the two modes ($a$ mode and $b$ mode) of the PCS, respectively. The various generation methods have been proposed theoretically in different physical systems (Agarwal, 1986; Gou et al., 1996; Zheng, 2002; Obada and Khalil, 2006; Dong et al., 2008; Gerry et al., 2011). The generation and characterization of PCS within spatially two confined superconducting cavities have been explored in recent experimental work (Gertler et al., 2023). As indicated in previous theoretical (Zheng, 2002) and experimental (Jeon et al., 2024) works, it also may possible to generate and manipulate of PCS in trapped ions systems. In order to facilitate the PCS in long-distance quantum information processing and communication tasks in most recent work (Truong et al., 2024) the experimentally feasible theoretical proposal to the generation of PCS in open space with high fidelity also introduced in detail. PCS are a type of two-mode non-Gaussian and nonclassical states (Hong and Guang-can, 1999), and its sub-Poissonian statistics, correlations in number fluctuations, squeezing, and violations of Cauchy–Schwarz inequalities (CSI) have been studied extensively (Agarwal, 1988; Lee, 1990; Obada et al., 2020; Duc et al., 2019; Truong et al., 2021). As non-Gaussian entangled states, PCS offers advantages in performing quantum tasks such as quantum communication (Tara and Agarwal, 1994; GÁBRIS and AGARWAL, 2007; Albert et al., 2019), interferometric lithography (Gong, 2013), quantum metrology (Giovannetti et al., 2006), quantum dialogue (Nguyen, 2004) and quantum key distribution (Zhou et al., 2010; Wang and Zhao, 2017).
@@
To study the effects of postselected von Neumann measurement on the PCS, we consider the PCS as the MD and its polarization as the measured system. As illustrated in Fig. 1, we simplify the setup by restricting the interaction between the MD and the measured system to the $a$-mode. The interaction Hamiltonian $H_{int}$ is of the von Neumann type, expressed as:
@@
$\times$ $|\psi_{i}\rangle\otimes|\phi\rangle,$
where $\Gamma=gt/\sigma$ represents the ratio between $gt$ and $\sigma$, $\mathbb{I}$ is the $2\times 2$ identity matrix, and $D(\Gamma/2)=e^{\frac{\Gamma}{2}(\hat{a}^{\dagger}-\hat{a})}$ is the displacement operator acting on the $a$ mode of the PCS. It is important to note that the coupling strength parameter $\Gamma$ characterizes the measurement strength in this study. The coupling between measured system and MD is referred to as weak when $\Gamma\text{$\ll 1$}$ and strong when $\Gamma\gg 1$. For this reason, $\Gamma$ can be considered the transaction parameter in our measurement model. The measurement process involves postselection. In this study, we assume the postselection of the system state to be $|\psi_{f}\rangle=|H\rangle$. After postselecting the system state, the final state of the MD is obtained, and its expression is given as
@@
with $P=\mathcal{N}_{\delta}^{2}e^{-\frac{\Gamma^{2}}{2}}\sum\limits_{n=0}^{\infty}\left[\frac{|\gamma|^{2n+\delta}}{n!(n+\delta)!}L_{n+\delta}^{(0)}\left(\Gamma^{2}\right)\right]$, and $\langle\sigma_{x}\rangle_{w}$ is the weak value of the system observable $\sigma_{x}$ defined as
@@
For simplify, we assume the weak value to be a real number by setting $\vartheta=0$. In calculating the normalization coefficient $\lambda$, the following expressions were used
@@
We have to mention that in the derivation of final state of MD, i.e., Eq. (10), we don’t use any approximation and the coupling strength parameter $\Gamma$ can take arbitrary value. If we interested in standard weak measurement regime, its enough to us only consider up to first-order term of the displacement operator $D\left(\frac{\Gamma}{2}\right)$ since $\Gamma\ll 1$ in this case, i.e., $D\left(\frac{\Gamma}{2}\right)\approx 1+$$\frac{\Gamma}{2}(a^{\dagger}-a)$. However, as studied in previous works (Turek et al., 2015a, b), the weak value $\langle\sigma_{x}\rangle_{w}$ in our scheme just a quantity of system observable $\sigma_{x}$ induced by particular pre- and post-selection processes on measured system, and it’s valid both in Aharonov’s weak measurement and von Neumann’s projective strong measurement regimes.
@@ III.1 Quadrature squeezing
As we can see, the values of $Q_{i}$ are bounded as $Q_{i}\geq-\frac{1}{4}$, and the $i$-th component of the quadrature operators of the PCS is said to be squeezed if $-\frac{1}{4}\leq Q_{i}<0$. After some algebra, we can obtain the specific expressions for the quadrature squeezing parameter $Q_{i}$ of the final MD state $|\Psi\rangle$, expressed as:
@@
To provide a clearer analysis of quadrature squeezing, we rely on numerical calculations, with the results presented in Figs. 2 and 3. In Fig. 2, we plot $Q_{1,\Psi}$ as a function of the parameter $\gamma$ for various coupling strength parameters $\Gamma$, while the weak value $\langle\sigma_{x}\rangle_{w}=5.671$ is fixed, corresponding to $\alpha=8\pi/9$. As shown in Fig. 2, $Q_{1,\phi}$ ($\Gamma=0$) remains positive and increases with $\gamma$ [see the thick green curve]. Additionally, Fig. 2 demonstrates that $Q_{1,\Psi}$ for cases where $\Gamma\neq 0$ also consistently takes positive values. This observation confirms that after the postselected von Neumann measurement, the $F_{1}$ quadrature of the final MD state $|\Psi\rangle$ still doesn’t exhibit squeezing, consistent with the behavior of the initial state.
@@
In Fig. 3, we present the numerical results for $Q_{2,\Psi}$, which corresponds to the quadrature $F_{2}$ of the final MD state $|\Psi\rangle$. In Fig. 3(a), $Q_{2,\Psi}$ is plotted as a function of the state parameter $\gamma$ for various coupling strength parameters $\Gamma$, with a fixed large weak value $\langle\sigma_{x}\rangle_{w}=5.671$, corresponding to $\alpha=8\pi/9$. Among these curves, the case $\Gamma=0$ (thick green curve) represents $Q_{2,\phi}$ for the initial state. As shown, $Q_{2,\phi}$ decreases with increasing $\gamma$ and converges to $-0.125$ across most parameter regions. Experimentally, the amount of squeezing is quantified in terms of decibels (dB) (Agarwal, 2013). This value of $-0.125$ corresponds to $9$ dB or $50\%$ squeezing. Notably, in the specific case where the PND is $\delta=0$, $Q_{2,\phi}$ is consistently less than zero. Thus, the initial PCS $|\phi\rangle$ with $\delta=0$ always exhibits squeezing along the $F_{2}$ quadrature.
@@
Figure 2: Quadrature squeezing $Q_{1,\Psi}$ as a function of the state parameter $\gamma$. The thick green line corresponds to the coupling strength parameter $\Gamma=0$, representing the case where the state is the initial PCS $|\phi\rangle$. The solid gray line represents $\Gamma=0.3$, while the purple, blue, and red dashed lines correspond to $\Gamma=0.5$, $\Gamma=0.7$, and $\Gamma=1$, respectively. For these calculations, we take $\delta=0$, $\alpha=\frac{8\pi}{9}$ and $\vartheta=0$
@@
In Fig. 3(a), after the postselected von Neumann measurements ($\Gamma\neq 0$), we observe that the value of $Q_{2,\Psi}$ is lower than $Q_{1,\phi}$ for the weak coupling strength parameter $\Gamma=0.3$. This behavior persists over a larger range of the state parameter $\gamma$, indicating that the squeezing characteristics of the quadrature $F_{2}$ are enhanced under postselected von Neumann measurements. However, as the coupling strength parameter $\Gamma$ increases, the value of $Q_{2,\Psi}$ decreases further compared to the $\Gamma=0.3$ case. For $\Gamma=0.3$, the minimum value of $Q_{2,\Psi}$ reaches approximately $-0.172$ near $\gamma=10$. This corresponds to a squeezing level of $11$dB or $69\%$. Therefore, after performing postselected WMs, the quadrature squeezing along the $F_{2}$ direction of the field increases by approximately $19$% under suitable system parameters and weak values of the measured system observable. To further investigate the signal amplification effect of weak values for weak system signals, Fig. 3(b) presents $Q_{2,\Psi}$ as a function of the weak value angle $\alpha$ for different coupling strength parameters $\Gamma$ and a fixed $\gamma=10$. The numerical results in Fig. 3(b) indicate that the weak value amplifies the squeezing effects of $Q_{2,\Psi}$ when the coupling strength parameter $\Gamma$ is small, highlighting the signal amplification capability of weak measurements. These results demonstrate that the squeezing effects of $Q_{2,\Psi}$ can be enhanced by postselected WMs.
@@
Figure 3: Quadrature squeezing $Q_{2,\Psi}$: (a) $Q_{2,\Psi}$ as a function of the state parameter $\gamma$ for different coupling strength parameters $\Gamma$, with the weak value fixed at $\langle\sigma_{x}\rangle_{w}=5.671$ corresponding to $\alpha=8\pi/9$. (b) $Q_{2,\Psi}$ as a function of the weak value parameter $\alpha$ for different coupling strength parameters $\Gamma$, with the state parameter $\gamma$ fixed at 10. Other parameters are the same as those in Fig. 2.